A forceacts between sodium and chloride ions of salt when put 1 cm apart in air The permittivity of air and dielectricconstant of water are E and K respectively when a piece of salt is put in water electrical force acting between sodium and chlorine ions 1 cm apart is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

when ions are placed in air

Fa= (1/4∏E) qq /(.01)2

when ions are placed in the water

Fw = (1/4∏EK) qq /(.01)2

      =Fa/K

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin